# Through the Lens of Split Vote: Exploring Disagreement, Difficulty and Calibration in Legal Case Outcome Classification

### 11 Feb 2024 v1

@@ 6.3.3 Results & Discussion
Tab 3 presents the experiment results. We make the following observations:
COC Performance COC Performance: Tab 3 indicates no improvement in COC performance (hm-F1) for SV cases in the test set, although improvement is observed for the development set (Tab 8 in App R). The scarcity of SV instances aside, temporal shifts in the dataset, as discussed in the longitudinal category (§ 5.1.2), may contribute to this discrepancy. SV cases could stabilize after a key legal decision. In other words, something that is legally uncertain in the training set time period may be settled law in the test dataset partitions. This suggests potential overfitting of our models to the training set timeframe. We encourage future models to incorporate the temporal dimension during training to mitigate such issues.
Confidence Calibration The implementation of TS greatly reduces ECE, suggesting enhanced confidence calibration for both models. However, fine-tuning with soft-loss to human labels yields minimal ECE changes with a discrepancy: BERT shows a slight decrease, while LegalBERT displays a minor increase. This mirrors the issue identified by Baan et al. 2022, highlighting the challenge of using ECE to measure calibration on disagreement data. They argue that even a classifier perfectly modeling human judgment distribution would still be severely miscalibrated when measured by ECE.
Human Calibration Soft-loss tuning slightly improves the alignment between the model’s prediction confidence and the human vote distribution (lower DistCE). Soft LegalBERT exhibits noticeably fewer instances of over-confident predictions when judges do not unanimously agree, as depicted on near-0/1 probability portions of Fig 3b, in contrast to the original model shown in Fig 3a. TS models exhibit substantially lower DistCE scores on split-vote cases than soft models. This may be attributed to an overly aggressive temperature, resulting in a more uniform output distribution (as shown in Fig 3c, with a temperature of 5.5). Moreover, judge votes often exhibit a quasi-bimodal distribution, with many split votes caused by a single judge dissenting pro/con the finding of a violation (refer to our discussion on single dissenting votes in § 4). Therefore, while TS models may display lower DistCE scores than soft models, they are not a good fit. It is crucial to analyze the distributions before drawing conclusion.
@@ 7 Conclusion
We present SV-ECHR, a COC dataset enriched with split-vote information. Our analysis of judges’ vote distributions leads to a taxonomy of disagreement sources. Our experiments reveal critical issues with current calibration measures, such as using confidence calibration metric ECE on data with inherent human disagreement. Additionally, caution is necessary when interpreting ECE and DistCE scores for human calibration. We call on the community to explore methods for measuring and improving the alignment of model calibration with human behavior; as well as more research into incorporating human disagreement and variability in NLP.
@@ Appendix E Correction of Inconsistent Metadata
Tab 6 shows the statistics about our correction of meta information. Notably, apart from the inconsistent allegation information pointed out by Xu et al., 2023b. We further found approximately 2% inconsistent violation information in the train set during our quality control process. Most of the cases with so-called striked-out cases which are difficult to parse. We also find finetuning on the corrected train set improves model’s COC performance (§ 6.1) as discussed in App M.
@@ Appendix G Debates in Legal Scholarship about Single Judge Dissenting Vote
Tab 5 reveals that $\sim$ 60% of chamber votes (involving 7 judges) exhibit entropy of around 0.4, indicating that one judge vote differently from the remaining 6 judges (entropy([1,6]) $\approx$ 0.4). A similar pattern is observed for Grand Chamber cases involving 17 judges. This Single Judge pattern aligns with ongoing debates in legal scholarship. For instance, Fobbe 2022 examined the number of dissenting opinions in decisions from the International Court of Justice (ICJ). Their results indicate a significant proportion of unanimous decisions, followed by a monotonously decreasing number of dissents. Some legal scholars support the view that diverging opinions do not signal a division of the bench inside the Court over the scope of protection of rights, even less a lowering of standards of protection. While other scholars, support the contrary view according to which dissenting opinions send a signal of walking-back in terms of effective protection. Tab 5 and Fobbe 2022’s finding actually shows that the case law does not consistently but rather seldomly give rise to dissenting opinions. Moreover, one can not ignore the National Judge’s dissenting vote (App I). It is difficult to assess what is the weight of a single judge dissenting on how the court will decide cases. Often times, dissenting opinions are not followed upon. What is of relevance is to determine a yardstick against which to measure disagreement in human decisions and this is what we are doing.
@@
Figure 6: Screenshot of Judges’ vote information in HUDOC
@@ Appendix I National Judge
Appendix M Ablation Experiment
@@
Appendix N Rationale for Avoiding Implementation of Uncertainty Evaluation Methods
@@
Appendix O Details of the Models
O.1 Architecture of the Hierarchical Model
@@
Appendix P Soft Loss Function
@@
Appendix Q Expected Calibration Error
@@
Appendix R More Calibration Results

### 16 Feb 2024 v2

@@ 6.3.3 Results & Discussion
Tab 3 confirms that SV cases are indeed the most challenging, reflected by the lowest COC performance (hm-F1). They also exhibit a higher degree of miscalibration (higher ECE) and misalignment with human responses (higher distCE). Applying TS greatly reduces ECE, without negatively impacting hm-F1, enhancing confidence calibration.
However, the models remain misaligned with the human vote distribution, as the high DistCE scores show and Fig 3 illustrates. Soft-loss tuning slightly improves the alignment between the model’s prediction confidence and the human vote distribution (lower DistCE). Soft LegalBERT exhibits noticeably fewer instances of over-confident predictions when judges do not unanimously agree, as depicted on near-0/1 probability portions of Fig 3b, in contrast to the original model shown in Fig 3a. TS models as expected reduce overly miscalibrated instances (drop in tail) but they exhibit substantially lower DistCE scores on split-vote cases than soft models. This may be attributed to an overly aggressive temperature, resulting in a more uniform output distribution (see Fig 3c, with a temperature of 5.5). Moreover, judge votes often exhibit a quasi-bimodal distribution, with many split votes caused by a single judge dissenting pro/con the finding of a violation (see discussion on single dissenting votes in § 4). Further, the DistCE score histograms (Fig 3d-f) illustrate that soft-tuning (Fig 3e) improves human alignment by reducing instances of extreme miscalibrations in the right tail, as compared to the original model in Fig 3d. It is noteworthy that the TS mode (Fig 3f) shifts the distribution further to the left by reducing the right tail. However, this is at the expense of predictions that perfectly align with human judgement probabilities, as evidenced by fewer instances of DistCE scores below 0.15 in the TS model compared to the Soft model. Therefore, despite TS models potentially displaying lower DistCE scores than soft models, they do not provide an optimal fit. It is crucial to analyze the distributions prior to drawing conclusions, especially as we have shown in cases with split votes, where distributions are bimodal.
@@ 7 Conclusion
We present SV-ECHR, a new COC dataset enriched with naturally occurring split-votes by judges on the ECtHR. We also present a SV-enriched taxonomy of disagreement sources. Our experiments reveal shortcomings when combining TS and ECE to improve and measure calibration of models against labels subject to inherent human disagreement. This is due to a distribution mismatch reflected in low DistCE scores. Soft loss training produces only slightly better human calibration scores. We call on the community to explore methods for measuring and improving the alignment of model calibration with human behavior; as well as more research into incorporating HLV in NLP.
@@ Appendix A DoF of Sources of Disagreement in Legal Decision Process
Figure 5: Screenshot of Judges’ vote information in HUDOC
@@ Appendix E Correction of Inconsistent Metadata
Tab 5 shows the statistics about our correction of meta information. Notably, apart from the inconsistent allegation information pointed out by Xu et al., 2023b. We further found approximately 2% inconsistent violation information in the train set during our quality control process. Most of the cases with so-called striked-out cases which are difficult to parse. We also find finetuning on the corrected train set improves model’s COC performance (§ 6.1) as discussed in App N.
@@ Appendix G Debates in Legal Scholarship about Single Judge Dissenting Vote
Tab 4 reveals that $\sim$ 60% of chamber votes (involving 7 judges) exhibit entropy of around 0.4, indicating that only one judge vote differently from the remaining 6 judges (entropy([1,6]) $\approx$ 0.4). A similar pattern is observed for Grand Chamber cases involving 17 judges. This Single Judge Dissenting pattern aligns with ongoing debates in legal scholarship. For instance, Fobbe 2022 examined the number of dissenting opinions in decisions from the International Court of Justice (ICJ). Their results indicate a significant proportion of unanimous decisions, followed by a monotonously decreasing number of dissents. Some legal scholars support the view that diverging opinions do not signal a division of the bench inside the Court over the scope of protection of rights, even less a lowering of standards of protection. While other scholars, support the contrary view according to which dissenting opinions send a signal of walking-back in terms of effective protection. Tab 4 and Fobbe 2022’s finding actually shows that the case law does not consistently but rather seldomly give rise to dissenting opinions. Moreover, one can not ignore the National Judge’s dissenting vote (App I). It is difficult to assess what is the weight of a single judge dissenting on how the court will decide cases. Often times, dissenting opinions are not followed upon. What is of relevance is to determine a yardstick against which to measure disagreement in human decisions and this is what we are doing.
@@ Appendix I National Judge
Appendix M The impact of domain-specific pre-training on uncertainty representations
Tab 6 reports the results of classification performance. Notably, the model with legal-specific pre-training (LegalBERT) outperforms the one with general pre-training (BERT).
Tab 8 shows that fine-tuning with soft-loss to human labels yields minimal ECE changes with a discrepancy: BERT shows a slight decrease, while LegalBERT displays a minor increase. This mirrors the issue identified by Baan et al. 2022, highlighting the challenge of using ECE to measure calibration on disagreement data. They argue that even a classifier perfectly modeling human judgment distribution would still be severely miscalibrated when measured by ECE.
mic-F1 mac-F1 hm-F1
BERT 72.31 ± 4.09 65.53 ± 6.22 64.22 ± 3.46
LegalBERT 73.55 ± 0.62 68.75 ± 1.13 68.03 ± 1.29
Table 6: COC performance on test set
Appendix N Ablation Experiment
@@
Appendix O Rationale for Avoiding Implementation of Uncertainty Evaluation Methods
@@
Appendix P Details of the Models
P.1 Architecture of the Hierarchical Model
@@
Appendix Q Soft Loss Function
@@
Appendix R Expected Calibration Error
@@
Appendix S More Calibration Results